40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2163     Entry time: Fri Oct 30 04:41:37 2009
Author: rob 
Type: Update 
Category: Locking 
Subject: working again 

I never actually figured out exactly what was wrong in entry 2162, but I managed to circumvent by changing the time sequence of events in the up script, moving the big gain increases in the common mode servo to the end of the script.  So the IFO can be locked again.

ELOG V3.1.3-